O’Neil Enterprises produces a line of canned soups for sale at supermarkets across the country. Demand has been “soft” recently and the company is operating at 70 percent of capacity. The company is considering dropping one of the soups, beef barley, in hopes of improving profitability. If beef barley is dropped, the revenue associated with it will be lost and the related variable costs saved. The CFO estimates that the fixed costs will also be reduced by 25 percent.

The following product line statements are available.

Product Broth Beef Barley Minestrone
Sales $ 38,100 $ 48,200 $ 56,600
Variable costs 23,800 42,200 43,700
Contribution margin $ 14,300 $ 6,000 $ 12,900
Fixed costs allocated to each product line 8,300 9,600 10,700
Operating profit (loss) $ 6,000 $ 3,600 $ 2,200

Required:

a-1. Complete the following differential cost schedule.

a-2. From an operating profit perspective, should O'Neil drop the beef barley line?

b. When the product manager for the minestrone soup hears that managers are considering dropping the beef barley line, she points out that many O’Neil customers buy more than one soup flavor and if beef barley is not available from O’Neil, some of them might stop buying the other soups as well. She estimates that 5 percent of the current sales of both broth and minestrone will be lost if beef barley is dropped.

b-1. Complete the following differential cost schedule.

b-2. Based on the estimate from the project manager, should O'Neil drop the beef barley line?
Answer #1
Differential Cost Schedule
Status Quo Alternative Difference
Revenue 142,900 94,700 -48,200
Less: VC 109,700 67,500 -42,200
Contribution Margin 33,200 27,200 -6,000
Less: Fixed Costs 28,600 21,450 -7,150
Operating Profit(Loss) 4,600 5,750 1,150
YES, since profit will increase by $1150
b-1
Differential Cost Schedule
Status Quo Alternative Difference
Revenue 142,900 89,965 -52,935
Less: VC 109,700 64,125 -45,575
Contribution Margin 33,200 25,840 -7,360
Less: Fixed Costs 28,600 21,450 -7,150
Operating Profit(Loss) 4,600 4,390 -210
b-2 No, as profit will reduce by $210

Ε в c 1 Differential Cost Schedule 2 Status Quo Alternative Difference 3 Revenue =38100+48200+56600 =38100+56600 =C3-B3 4 Les